### v2

# TFusionOcc: T-Primitive Based Object-Centric Multi-Sensor Fusion Framework for 3D Occupancy Prediction

Abstract
The prediction of 3D semantic occupancy enables autonomous vehicles (AVs) to perceive the fine-grained geometric and semantic scene structure for safe navigation and decision-making. Existing methods mainly rely on either voxel-based representations, which incur redundant computation over empty regions, or on object-centric Gaussian primitives, which are limited in modeling complex, non-convex, and asymmetric structures. In this paper, we present TFusionOcc, a T-primitive-based object-centric multi-sensor fusion framework for 3D semantic occupancy prediction. Specifically, we introduce a family of Student’s t-distribution-based T-primitives, including the plain T-primitive, T-Superquadric and deformable T-Superquadric with inverse warping, where the deformable T-Superquadric serves as the key geometry-enhancing primitive. We further develop a unified probabilistic formulation based on the Student’s t-distribution and the T-mixture model (TMM) to jointly model occupancy and semantics, and design a tightly coupled multi-stage fusion architecture to effectively integrate camera and LiDAR cues. Extensive experiments on nuScenes show state-of-the-art performance, while additional evaluations on nuScenes-C demonstrate strong robustness under most corruption scenarios. The code will be available at: https://github.com/DanielMing123/TFusionOcc
Index Terms:
Autonomous driving, 3D semantic occupancy prediction, Multi-sensor fusion, Environment perception
I Introduction
Accurate and efficient scene representation is essential for autonomous driving, since downstream planning and decision-making rely on a reliable understanding of both geometry and semantics. Compared with conventional 3D object detection and BEV segmentation, 3D semantic occupancy prediction provides a more complete and denser description of the surrounding environment, especially for irregular structures, free space, and partially occluded regions. Multi-sensor fusion further improves this representation by combining the complementary strengths of cameras and LiDAR.
Figure 1: Pipeline of three approaches: Voxel-based approach (top), 3D-Gaussian-Primitive-based object-centric approach (middle), and our approach (bottom).
Recent progress in multi-sensor 3D semantic occupancy prediction has been largely driven by voxel-based fusion methods [18, 21, 37, 16, 35, 4, 26], which integrate complementary signals from cameras and LiDAR. Although these methods have achieved strong performance, they generally rely on dense volumetric processing (Fig. 1, voxel-based approach), leading to substantial redundant computation over empty regions that dominate outdoor driving scenes. This inefficiency has motivated the development of object-centric representations, where a compact set of learnable primitives is used to model occupied regions more selectively. Among object-centric approaches[39, 23], Gaussian-based primitives have recently shown promise for improving efficiency and compact representation (Fig 1, 3D Gaussian-based approach). However, their geometric modeling capability remains limited when faced with non-convex, asymmetric, and structurally irregular scene elements that are common in real driving environments. In addition, Gaussian-like formulations are less effective in handling outliers and long-tailed spatial distributions, which may reduce robustness under noisy or corrupted observations. As a result, existing object-centric occupancy frameworks still face a performance gap when compared with strong voxel-based methods, especially when an accurate geometric representation is required.
To address these limitations, we propose TFusionOcc (Fig 1, our approach), a T-primitive-based object-centric multi-sensor fusion framework for 3D semantic occupancy prediction. Specifically, we introduce a family of Student’s t-distribution-based T-primitives, including the plain T-primitive, T-Superquadric and a deformable T-Superquadric with inverse warping, where the latter serves as the key geometry-enhancing primitive. We further develop a unified probabilistic formulation based on Student’s t-distribution and the T-mixture model (TMM) to jointly model occupancy and semantics, and design a compact multi-stage fusion pipeline to effectively integrate camera and LiDAR cues. In this way, TFusionOcc improves geometric representation while preserving the efficiency advantage of object-centric modeling. Extensive experiments on nuScenes [2] show that TFusionOcc achieves state-of-the-art (SOTA) overall performance and consistently benefits from stronger primitive geometry. In particular, our results indicate that the flexibility of primitive shapes is more decisive than simply increasing the primitive count, highlighting the importance of geometry-aware primitive design. Additional evaluations on nuScenes-C [32] in different corruption settings further demonstrate strong robustness in many challenging cases. The main contributions of this paper are summarized as follows:
  • We propose a family of Student’s t-distribution-based T-primitives for object-centric 3D occupancy prediction, including the plain T-primitive, T-Superquadric, and a deformable T-Superquadric with inverse warping. In particular, the deformable T-Superquadric significantly improves geometric modeling flexibility, enabling better representation of non-convex and asymmetric structures.
  • We develop a unified probabilistic modeling framework based on Student’s t-distribution and the T-mixture model, through which both occupancy and semantics are jointly modeled. This design improves robustness to outliers and complex real-world noise while alleviating the limitations of Gaussian-based primitives in modeling long-tailed spatial distributions.
  • We design a tightly coupled multi-sensor fusion pipeline to support effective T-primitive learning. Specifically, it includes fused-depth-guided visual feature lifting, transformer-based T-primitive-centric refinement, and the GCWSFusion module for adaptive cross-modal interaction, allowing T-primitives to better exploit cross-modal cues for 3D semantic occupancy prediction.
The remainder of this paper is structured as follows. Section II provides an overview of related research and identifies the key differences between this study and previous publications. Section III outlines the general framework of TFusionOcc and offers a detailed explanation of the implementation of each module. Section IV presents the results of our experiments. Finally, Section V provides the conclusion of our work.
II Related Work
II-A Voxel-Based Approach
In OccFusion [18], the authors extract a set of 3D voxel volumes from each different modality sensor, such as surround-view cameras, surround-view radars and LiDAR, followed by the feature fusion operation through the proposed dynamic feature fusion module to perform multi-sensor fusion-based 3D semantic occupancy prediction. The model exhibits strong robustness under challenging rain and night scenarios. In Co-Occ [21], the authors leverage the volume rendering technique, which was originally derived from NeRF [15], to refine the intermediate 3D feature volumes. They also proposed a GSFusion module for efficient multimodal feature fusion, achieving SOTA performance. In FusionOcc [37], the authors proposed a cross-modality fusion module to generate a high-quality depth-aware visual feature, followed by a 2D to 3D view transformation, resulting in volumes of depth-aware visual features under the Cartesian coordinate system. The LiDAR branch conducts voxelization and voxel encoding, resulting in a LiDAR feature volume also under the Cartesian coordinate system. Different modality feature volumes were fused via feature channel concatenation and further refined by a 3D Encoder. In OccCylindrical [16], the authors proposed a multi-sensor fusion-based framework that uses a cylindrical partition to perform voxelisation of a LiDAR-captured 3D point cloud and a pseudo 3D point cloud predicted by surround-view cameras, resulting in better preservation of 3D geometric information. The feature fusion and further refinement operations are performed under the cylindrical coordinate system. In AdaptiveOcc [34], the author explored using an efficient octree data structure to preserve the fine-grained geometry of the 3D scene, achieving a good balance between computational complexity and prediction performance for the voxel-grid-based approach. In DAOcc [35], the authors leverage multi-task learning to introduce an additional 3D supervision signal to guide intermediate features, leading to robust generalization and overall SOTA performance. Similarly, in Doracamom [40], the authors developed a joint 3D detection and occupancy prediction pipeline under a camera-radar fusion paradigm. By leveraging the strengths of multi-task learning and further incorporating temporal fusion, the proposed method achieved superior performance in both 3D object detection and 3D semantic occupancy prediction. In addition, SDGOcc [4] incorporates a multi-task head to introduce a 2D semantic mask into the model training procedure, leveraging an additional 2D semantic mask to enhance the precision and densification of depth maps obtained from LiDAR point cloud projection, thereby boosting the model’s overall performance.
Despite achieving remarkable performance, these voxel-based methods incur redundant calculations for empty voxels, which are often dominant in 3D driving scenarios, resulting in inefficient computation. Our study presents a novel object-centric approach, aiming to address the aforementioned constraints by leveraging T-Primitives to model only occupied space in driving scenarios.
II-B Object-Centric Modeling Approach
In the GaussianFormer series [9, 7], the authors proposed an object-centric 3D Gaussian model together with a Gaussian mixture model (GMM) to improve computational efficiency and robustness. In GraphGSOcc [27], the authors combine the object-centric 3D Gaussian representation with a graph transformer to realize dynamic and static object decoupling modeling. Allow the model to optimize foreground dynamic objects and background static objects separately. The approach further incorporates a temporal fusion mechanism to achieve superior performance. In GaussianFormer3D [39], the authors extend the GaussianFormer work to a multi-sensor fusion scenario. Using a set of 3D Gaussian primitives as an intermediate feature representation, the authors designed a pipeline that enables each 3D Gaussian primitive to aggregate features from different modalities and fuse them to perform 3D primitive to-3D voxel volume Gaussian splatting. Benefiting from the object-centric modeling technique and multi-modal fusion, the model achieves significant performance and efficiency improvements. In GaussianFusionOcc [23], the authors proposed a GaussianFusionBlock module that performs efficient and high-quality feature aggregation and fusion across different modalities, allowing high-quality 3D Gaussian primitive modeling and leading to outstanding performance.
Unlike previous object-centric methods, our work introduces a family of Student’s t-distribution-based T-primitives for 3D semantic occupancy prediction. In particular, we propose a deformable T-Superquadric with inverse warping, which substantially strengthens primitive-level geometric flexibility. We further formulate occupancy and semantics under a unified TMM framework and design a multi-stage camera–LiDAR fusion pipeline to support effective T-primitive refinement. In this way, our method aims to preserve the efficiency advantage of object-centric modeling while improving geometric representation capability and robustness.
III TFusionOcc
This work leverages a set of Student-T-distribution-based primitives combined with a 3D to 3D splatting operation to generate a dense 3D semantic occupancy grid of the surrounding scene by integrating information from surround-view cameras and LiDAR. Thus, the problem can be formulated as follows:
$P^{T}_{refined}=F\left(Cam^{1},Cam^{2},...,Cam^{N},Lidar,P^{T}\right)$ (1)
$Occ=Splat\left(P^{T}_{refined}\right)$ (2)
where $F$ refers to a multi-sensor fusion framework, $P^{T}$ refers to a set of primitives based on T-distribution $P^{T}=\left\{p^{T}_{1},p^{T}_{2},...,p^{T}_{K}\right\}$, along with a set of query vectors $Q=\left\{q_{1},q_{2},...,q_{K}\right\}$, $P^{T}_{refined}$ refers to a set of T-distribution-based primitives refined through multi-sensor fusion framework, $N$ refers to the total number of surround-view images and $Splat\left(\cdot\right)$ refers to the splatting operation. In the final predicted 3D semantic occupancy $Occ\in R^{\left\{X\times Y\times Z\right\}}$, each grid is assigned a semantic property ranging from 0 to $C$, where $C$ refers to the total number of semantic classes. In our case, a class value of 0 corresponds to an empty grid.
III-A Overview
Figure 2: Overall architecture of TFusionOcc. The pipeline comprises a camera branch, a LiDAR branch, and a multi-stage fusion branch. The camera branch extracts multi-scale visual features and predicts a pseudo 3D point cloud from surround-view images. The pseudo 3D point cloud is further projected and cylindrically partitioned, resulting in camera-based, multi-scale, dense depth maps and a voxel volume defined under cylindrical coordinates. The LiDAR branch applies a cylindrical partition followed by a 3D encoder to extract the LiDAR feature. Meanwhile, the LiDAR point cloud is also projected to generate LiDAR-based, multi-scale, sparse depth maps. The feature fusion branch adopts a multi-stage fusion strategy to merge all outputs from the two-modality branch and leverages a proposed transformer to refine the T-primitives property through fused features.
The overall architecture is shown in Fig. 2. For the camera branch, given $\left\{img^{1},img^{2},...,img^{N}\right\}$ surround-view images, a 2D encoder and a decoder are leveraged to extract $V^{Cam}=\left\{\left\{V_{n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$ multi-scale visual features. Meanwhile, a DepthNet is leveraged, resulting in a camera-based pseudo 3D point cloud $Pt^{cam}$ and a multi-scale camera-based dense depth maps $V_{depth}^{cam}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$, where $D$ refers to the total amount of depth-bins. The pseudo 3D point cloud $Pt^{cam}$ is further divided into cylindrical voxels, resulting in a camera-based 3D voxel volume $Vol^{Cam}_{Cylin}$ defined in cylindrical coordinates. For the LiDAR branch, given a 3D point cloud $Pt^{lidar}$ from LiDAR, the cylindrical partition is first applied, resulting in a LiDAR-based 3D voxel volume $Vol^{Lidar}_{Cylin}$. Then a 3D encoder (e.g., Cylinder3D [41]) is used to extract LiDAR features $V^{Lidar}$ in cylindrical coordinates. Meanwhile, $Pt^{Lidar}$ is projected onto each surround-view camera frame, resulting in sparse multi-scale LiDAR-based depth maps $V_{depth}^{lidar}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$. Then, a multi-stage feature fusion strategy is adopted. For the camera branch side, $V_{depth}^{cam}\oplus V_{depth}^{lidar}$ results in multi-scale fused dense depth maps $V_{depth}^{fuse}$. Then $V_{depth}^{fuse}$ is encoded by a multi-layer perceptron (MLP) and added to $V^{Cam}$, resulting in multi-scale depth-aware visual features $F_{depth}^{Cam}$. In addition, an outer product operation $F_{depth}^{Cam}\otimes V_{depth}^{fuse}$ is conducted, resulting in a multi-scale lifted depth-aware visual feature $Vol^{Cam}_{depth}$. For the LiDAR branch side, each occupied voxel center of $Vol^{Lidar}_{Cylin}$ serves as an anchor to project onto $V^{Cam}$, thereby aggregating semantic information. The aggregated semantic information is further fused with the LiDAR feature $Vol^{Lidar}_{Cylin}$ through an early-fusion module, resulting in a semantic-aware LiDAR feature $Vol^{Lidar}_{sem}$. The $Vol^{Cam}_{Cylin}$ and $Vol^{Lidar}_{Cylin}$ are fused through the skeleton merge module, resulting in a fused voxel volume $Vol^{Fused}_{Cylin}$ under cylindrical coordinates. The $Vol^{Fused}_{Cylin}$ is then used as primitive anchors to initialize T-Primitives $P^{T}$ under cylindrical coordinates. The $P^{T}$, $Vol^{Cam}_{depth}$ and $Vol^{Lidar}_{sem}$ are fed to the transformer module, which stacks the $N\times$ blocks to obtain a set of refined T-Primitives. Lastly, the refined set of T-Primitives performs a splatting operation, resulting in the final 3D semantic occupancy grid.
III-B Camera-based Pseudo 3D Point Cloud and Multi-Scale Cam Dense Depth Maps Generation
Figure 3: Inner Structure of DepthNet. The 1/8-scale visual features are first used to generate a 1/8-scale depth map. Then, bilinear interpolation is leveraged to generate 1/16- and 1/32-scale depth maps. Meanwhile, the image-based pseudo-point cloud is generated solely from a 1/8-scale depth map.
The pseudo 3D point cloud $Pt^{cam}$ and camera-based multi-scale dense depth maps $V^{cam}_{depth}$ are generated through the DepthNet module as shown in Fig. 3. The DepthNet module takes $V^{Cam}_{Encode}$ as input and leverages a SECONDFPN [33] to generate a single-scale visual feature $V^{cam}_{sec}=\left\{V_{n}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}$, whose resolution is $\frac{1}{8}$ of the input image. The $V^{cam}_{sec}$ is further passed to an MLP to perform the multi-hot per-ray depth estimation, resulting in a depth map $V_{depth\_1}\in\mathbb{R}^{N\times D\times H_{l}\times W_{l}}$ with each pixel ray having a set of depth bins separated by 0.5m intervals, and each depth bin on the pixel ray has a probability range in $[0,1]$. Following [7], we adopt the 3D semantic occupancy label to supervise $V_{depth\_1}$. Then, bilinear interpolation with a 0.5 downsampling factor is applied iteratively to $V_{depth\_1}$, resulting in two additional depth maps that have $\frac{1}{16}$ and $\frac{1}{32}$ resolution of the input image. The predicted depth map and the two extra downsampled depth maps form $V_{depth}^{cam}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$. Meanwhile, the pseudo 3D point cloud $Pt^{cam}$ is acquired only from $\frac{1}{8}$ scale depth map $V_{depth\_1}$.
III-C Multi-scale-Fused Dense Depth-Maps Generation
Figure 4: Multi-Scale Fused Dense Depth Maps Generation. This demonstrates the detailed multi-source depth map fusion from a per-pixel ray perspective.
Given the multi-scale dense depth maps from cameras, $V^{cam}_{depth}$, and the multi-scale sparse depth maps from LiDAR, $V^{lidar}_{depth}$, we perform element-wise addition at the per-pixel level, followed by a value clamping operation. This process produces the final multi-scale fused dense depth maps, as illustrated in Fig. 4. All depth maps from different sensors share a consistent representation, with the same number of depth bins $D$ and a uniform interval of 0.5 m between adjacent bins along each pixel ray. However, their value distributions differ: on the camera side, $V^{cam}_{depth}$ is represented as multi-hot per-pixel ray depth maps, where each depth bin takes a value within the range $[0,1]$. In contrast, $V^{lidar}_{depth}$ is represented as single-hot per-pixel ray depth maps, where only one depth bin is assigned a value of 1.0. When the camera and LiDAR depth maps are combined via per-pixel addition, the resulting fused pixel rays may contain depth-bin values exceeding 1.0. To ensure valid probability values, all depth-bin responses are subsequently clamped to the range $[0,1]$.
This fusion strategy combines the complementary strengths of the two modalities. The dense camera depth maps provide rich structural coverage but are less accurate, whereas the LiDAR depth maps are sparse but geometrically precise. Their combination yields denser, more reliable fused depth maps, which are subsequently used to enhance visual features and facilitate cross-modal fusion.
III-D Skeleton Merge Module
Figure 5: Skeleton Merge Module. The upper LiDAR-branch serves as the main skeleton to provide a foundation structure for the 3D scene, and the bottom camera-branch serves as an augmentation based on the main skeleton to provide more detailed local structure to compensate for the fine-grained geometry of the main skeleton.
The primitive anchor initialization process is shown in Fig. 5. Given the pseudo 3D point cloud derived from the camera $Pt^{cam}$ and the real 3D point cloud of the LiDAR $Pt^{lidar}$, we first apply the cylindrical partition to both point clouds, resulting in two volumes of voxels $Vol^{Cam}_{Cylin}$ and $Vol^{Lidar}_{Cylin}$ under cylindrical coordinates. We then predefine the total primitive anchors of $M+N$, where the anchors $M$ are sampled from $Vol^{Lidar}_{Cylin}$ and the anchors $N$ are sampled from $Vol^{Cam}_{Cylin}$, with a ratio of $M:N=3:1$. We adopt the farthest point sampling strategy to obtain $M$ LiDAR anchors from $Vol^{Lidar}_{Cylin}$, which act as the main structural skeleton. Using these LiDAR anchors as reference, any voxel in $Vol^{Cam}_{Cylin}$ that spatially overlaps with the LiDAR anchors is removed. In addition, we impose a radius constraint r=5r=5m to perform range filtering. Any voxel in $Vol^{Cam}_{Cylin}$ whose closest LiDAR anchor is farther away than $r$ is also discarded. Then, $N$ Camera anchors are obtained using the same sampling strategy on the remaining voxels in $Vol^{Cam}_{Cylin}$. Based on the aforementioned filtering and sampling strategy, we acquire a final set of $M+N$ anchors whose locations will be used to initialize a set of T-primitives.
This design exploits the complementary strengths of the two modalities: LiDAR provides the main geometric structure with higher spatial accuracy, while the camera branch supplements local details and occluded regions. The resulting fused skeleton, therefore, offers a better initialization for subsequent T-primitive modeling.
III-E Early-Fusion Module
Figure 6: Early-Fusion Module. Each occupied voxel center in $Vol^{Lidar}_{Cylin}$ serves as an anchor and is projected onto multi-scale visual features to aggregate semantic information, yielding a semantic-aware LiDAR feature.
The detailed inner structure of the Early-Fusion module is shown in Fig. 6. Given the sparse volume of the voxel $Vol_{Cylin}^{Lidar}$ in cylindrical coordinates, we leverage each occupied voxel center as an anchor and convert each anchor position from cylindrical coordinates to Cartesian coordinates. Then, each anchor is projected onto multi-scale visual features $V^{Cam}$ to aggregate semantic information at three scales. The aggregated three-scale visual features are weighted, summed, and concatenated back to the corresponding voxel-associated LiDAR feature, then fed through an MLP layer to perform the feature fusion, resulting in the last semantic-aware LiDAR feature $Vol_{sem}^{Lidar}$.
This module injects image semantics into sparse LiDAR features at an early stage, allowing the LiDAR branch to retain its geometric precision while gaining richer semantic context from the camera branch.
III-F T-Primitive Representation
We propose a probabilistic superposition of Student’s t-distribution-based primitives as an efficient and effective 3D scene representation. Specifically, we propose three types of primitives that are enhanced by the inner kernel of the T-distribution or T-distribution-like. The first is the Student-T-distribution-based plain primitive, which has the attributes as follows:
$P_{i}^{T}(m_{i},s_{i},r_{i},\alpha_{i},c_{i})$ (3)
where $m_{i}$, $s_{i}$, $r_{i}$, $\alpha_{i}$ and $c_{i}$ refer to the mean, scale value for the $x,y,z$ axes, quaternion vector, opacity and semantic vector of the $i$-th primitive, respectively. The second is the Student-T-distribution-based superquadric primitive, which has the attributes as follows:
$P_{SQi}^{T}(m_{i},s_{i},r_{i},\alpha_{i},\epsilon_{1_{i}},\epsilon_{2_{i}},c_{i})$ (4)
where $\epsilon_{1}$ and $\epsilon_{2}$ are the shape exponents of the superquadric. The third is the Student-T-distribution-based superquadric with inverse-warp primitive that has the attributes as follows:
$P_{SQIWi}^{T}(m_{i},s_{i},r_{i},\alpha_{i},\epsilon_{1_{i}},\epsilon_{2_{i}},\left\{\omega_{i,k}\right\}_{k=1}^{24},c_{i})$ (5)
where $\left\{\omega_{i,k}\right\}_{k=1}^{24}$ is a set of learnable weighting scalers used to determine the magnitude of applied deformation, and each $\omega_{i,k}$ has a value range between -1 and 1.
Inspired by [7], we perform 3D geometry prediction and semantic prediction separately and adopt the multiplication theorem of probability and the T-mixture model to address them probabilistically.
III-F1 Student-T-Distribution and Multiplication Theorem Based Geometry Prediction
In practice, we interpret the T-distribution primitives $P^{T}=\left\{P^{T}_{i}\right\}_{i=1}^{K}$ as the probability of their surrounding space being occupied. In detail, we assign a probability value of 100% to the centers of the $P^{T}_{i}$, which decays with respect to the distance from the centers $m_{i}$. Assuming conditional independence of occupancy among different T-primitives, the final occupancy probability at x can be computed as:
$p_{o}\left(x\right)=1-\prod_{i=1}^{K}\left(1-p\left(x|P^{T}_{i}\right)\right)$ (6)
where $p\left(x|P^{T}_{i}\right)$ denotes the probability of point x being occupied induced by the $i$-th T-primitive. Depending on the primitive type, it is instantiated by (8), (12), or (16). $p_{o}\left(x\right)$ refers to the overall probability of occupancy at point x.
III-F2 T-Mixture Model Based Semantic Prediction
A set of T-primitives can be considered as a T-mixture model, and the final semantic predictions can be calculated via a weighted aggregation of semantic probabilities from all contributing T-primitives in the scene as:
$p_{sem}\left(x\right)=\frac{{\textstyle\sum_{i=1}^{K}p\left(x|P^{T}_{i}\right)a_{i}\tilde{c}_{i}}}{{\textstyle\sum_{j=1}^{K}p\left(x|P^{T}_{j}\right)a_{j}}}$ (7)
where $\tilde{c}_{i}$ represents the softmax-normalized semantic property of $i$-th T-primitive, and $a_{i}$ refers to $l1$ normalized opacity property of $i$-th T-primitive.
III-F3 Student-T-Distribution Based Plain Primitive
The plain T-distribution primitive $p\left(x|P^{T}_{i}\right)$ with the lowest geometric flexibility can be formulated as:
$p\left(x|P^{T}_{i}\right)$ $=\frac{\Gamma\left(\frac{\nu_{i}+3}{2}\right)}{\Gamma\left(\frac{\nu_{i}}{2}\right)\cdot\left(\nu_{i}\pi\right)^{3/2}\cdot\left|\Sigma\right|^{1/2}}\cdot$
$\left[1+\frac{1}{\nu_{i}}\left(x-m\right)^{T}\Sigma^{-1}\left(x-m\right)\right]^{-\frac{\nu_{i}+3}{2}}$ (8)
where $\nu_{i}$ denotes the degrees of freedom of the $i$-th T-primitive. In our setting, we treat the number of voxels intersected by a T-primitive in the final 3D voxel grid as an approximation of its observable local support. Therefore, we adaptively define
$\nu_{i}=N_{i}-1,$ (9)
where $N_{i}$ is the number of intersected voxels associated with the $i$-th T-primitive. In this way, primitives supported by fewer voxels retain heavier tails for better robustness, while primitives supported by more voxels yield more concentrated distributions. This adaptive design allows each T-primitive to adjust its probabilistic shape according to its local geometric support in the scene. $\Gamma\left(\cdot\right)$ denotes the gamma function, which satisfies $\Gamma\left(n\right)=\left(n-1\right)!$ for positive integers $n$. Compared to the plain Gaussian-primitive, this T-primitive is highly tailed in its probability distribution, leading to better robustness against outliers.
III-F4 Student-T-Distribution Based SuperQuadric Primitive
The T-probabilistic modeling mechanism is adopted to convert superquadrics into occupancy probabilities. We first transform the 3D point position $(x,y,z)$, which is occupied by the $i$-th T-Superquadric primitive, into its local coordinate system as:
$\begin{bmatrix}X_{S_{i}}\\ Y_{S_{i}}\\ Z_{S_{i}}\\ \end{bmatrix}=R_{i}\times\begin{bmatrix}x-m_{x_{i}}\\ y-m_{y_{i}}\\ z-m_{z_{i}}\\ \end{bmatrix}$ (10)
where $(X_{S_{i}},Y_{S_{i}},Z_{S_{i}})$ refers to the local coordinate under the $i$-th T-Superquadric primitive, and $R_{i}$ is the rotation matrix of the $i$-th T-Superquadric primitive. Then, the occupancy probability of the 3D point $(x,y,z)$ associated with the $i$-th T-Superquadric primitive can be computed as:
$f\left(X_{S_{i}},Y_{S_{i}},Z_{S_{i}}\right)=\left(\left(\frac{X_{S_{i}}}{S_{x_{i}}}\right)^{\frac{2}{\epsilon_{2_{i}}}}+\left(\frac{Y_{S_{i}}}{S_{y_{i}}}\right)^{\frac{2}{\epsilon_{2_{i}}}}\right)^{\frac{\epsilon_{2_{i}}}{\epsilon_{1_{i}}}}+\left(\frac{Z_{S_{i}}}{S_{z_{i}}}\right)^{\frac{2}{\epsilon_{1_{i}}}}$ (11)
$p\left(x|P^{T}_{SQi}\right)=\left(1+\frac{1}{\nu_{i}}\cdot f\left(X_{S_{i}},Y_{S_{i}},Z_{S_{i}}\right)\right)^{-\frac{\nu_{i}+3}{2}}$ (12)
where $(S_{x_{i}},S_{y_{i}},S_{z_{i}})$ refers to the scale parameters of $i$-th T-Superquadric primitive along the $X,Y,Z$ axes and $\nu_{i}$ refers to the degree of freedom, and has the same definition as in (8).
III-F5 Student-T-Distribution Based SuperQuadric With Inverse-Warp Primitive
So far, thanks to the T-distribution kernel enhancement, the T-Superquadric primitive is robust to outliers. Nevertheless, due to limited shape variety, it cannot capture non-convex, asymmetric shapes, which are very common in the daily 3D driving scene. To solve this limitation, we make the T-Superquadric primitive deformable through inverse-warp enhancement. Specifically, we choose 24 basis field functions $\left\{B_{k}(u,v,w)\right\}_{k=1}^{24}$ to deform the local coordinates $(X_{S_{i}},Y_{S_{i}},Z_{S_{i}})$ as follows:
$u=\frac{X_{S_{i}}}{S_{x_{i}}},v=\frac{Y_{S_{i}}}{S_{y_{i}}},w=\frac{Z_{S_{i}}}{S_{z_{i}}}$ (13)
$\begin{bmatrix}\tilde{X}_{S_{i}}\\ \tilde{Y}_{S_{i}}\\ \tilde{Z}_{S_{i}}\\ \end{bmatrix}=\begin{bmatrix}X_{S_{i}}\\ Y_{S_{i}}\\ Z_{S_{i}}\\ \end{bmatrix}-\sum_{k=1}^{24}\omega_{i,k}\cdot B_{k}\left(u,v,w\right)$ (14)
Then, the occupancy probability of the 3D point associated with $i$-th T-Superquadric-Inverse-Warp primitive can be calculated as follows:
$f\left(\tilde{X}_{S_{i}},\tilde{Y}_{S_{i}},\tilde{Z}_{S_{i}}\right)=\left(\left(\frac{\tilde{X}_{S_{i}}}{S_{x_{i}}}\right)^{\frac{2}{\epsilon_{2_{i}}}}+\left(\frac{\tilde{Y}_{S_{i}}}{S_{y_{i}}}\right)^{\frac{2}{\epsilon_{2_{i}}}}\right)^{\frac{\epsilon_{2_{i}}}{\epsilon_{1_{i}}}}+\left(\frac{\tilde{Z}_{S_{i}}}{S_{z_{i}}}\right)^{\frac{2}{\epsilon_{1_{i}}}}$ (15)
$p\left(x|P^{T}_{SQIWi}\right)=\left(1+\frac{1}{\nu_{i}}\cdot f\left(\tilde{X}_{S_{i}},\tilde{Y}_{S_{i}},\tilde{Z}_{S_{i}}\right)\right)^{-\frac{\nu_{i}+3}{2}}$ (16)
Basis Field $u$ $v$ $w$ $B_{1}$ 1.0 0.0 0.0 $B_{2}$ 0.0 1.0 0.0 $B_{3}$ 0.0 0.0 1.0 $B_{4}$ $u$ 0.0 0.0 $B_{5}$ 0.0 $v$ 0.0 $B_{6}$ 0.0 0.0 $w$ $B_{7}$ $v$ 0.0 0.0 $B_{8}$ $w$ 0.0 0.0 $B_{9}$ 0.0 $w$ 0.0 $B_{10}$ 0.0 $u$ 0.0 $B_{11}$ 0.0 0.0 $u$ $B_{12}$ 0.0 0.0 $v$ $B_{13}$ wv-w*v wuw*u 0.0 $B_{14}$ 0.0 uw-u*w uvu*v $B_{15}$ vwv*w 0.0 vu-v*u $B_{16}$ $w^{2}$ 0.0 0.0 $B_{17}$ 0.0 $w^{2}$ 0.0 $B_{18}$ 0.0 0.0 $u^{2}+v^{2}$ $B_{19}$ $u^{2}$ 0.0 0.0 $B_{20}$ 0.0 $v^{2}$ 0.0 $B_{21}$ 0.0 0.0 $w^{2}$ $B_{22}$ $(u^{2}+v^{2})*u$ $(u^{2}+v^{2})*v$ 0.0 $B_{23}$ uvu*v uvu*v 0.0 $B_{24}$ $u*v^{2}$ $u^{2}*v$ 0.0
TABLE I: The 24 basis field functions make the T-Superquadric deformable. Those functions enable full shears, twists, bends, quadratics, radial bulge, and smooth corners.
Each basis field function is listed in Table I. $B_{1}$ to $B_{9}$ provide constants, linear, basic shears, and taper deformation. $B_{10}$ to $B_{18}$ enable full shear, twist, and bending deformation. Moreover, $B_{19}$ to $B_{24}$ provide quadratics, radial bulge, and smooth corner deformation.
III-G Transformer
III-G1 Fused Depth Map Guided 3D Deformable Attention
Figure 7: Fused Depth Map Guided 3D Deformable Attention For the same visual feature at location $(u_{i},v_{j})$, several different depth values ranging between $[0.0,1.0]$ along the depth-axis are used to weight the same pixel feature, resulting in the depth encoded visual feature to mitigate the projection ambiguity problem stemming from the projection process.
Taking multi-scale lifted depth-aware visual features and a set of T-primitive as input, we use a 3D deformable attention operator, namely 3DFA [11], which performs feature aggregation on lifted visual features to mitigate the 3D to 2D projection ambiguity problem, to aggregate visual features for each T-primitive. As shown in Fig. 2 camera branch and Fig. 7, the multi-scale lifted depth-aware visual features $Vol^{Cam}_{depth}$ are obtained through outer product operation $F_{depth}^{Cam}\otimes V_{depth}^{fuse}$, this operation gives an accurate depth encoding for each visual feature in the feature map. For each T-primitive feature sampling $p^{T}_{i}$, we first use the mean location $m_{i}$ of each $p^{T}_{i}$ as the center, convert the center location from the cylindrical coordinate to the cartesian coordinate, and feed its associate query feature $q_{i}$ to an MLP layer, resulting in 13 offsets $\Delta m_{j}$ with respect to the center to form a set of 3D sampling points $m^{samp}=\left\{m^{samp}_{i}=m_{i}+\Delta m_{j}\mid j=1,...,13\right\}$. Then, we leverage extrinsic and intrinsic matrices to project each 3D sampling point $m^{samp}_{i}$ onto the lifted depth-aware visual features, resulting in the location of the sampling of the pixel frame $\bar{m}^{samp}_{i}=\left(u_{i},v_{i},d_{i}\right)$. Lastly, a 3D deformable attention operator 3DFA takes $\bar{m}^{samp}_{i}$ as a sampling location to perform cross-attention visual feature aggregation on $Vol^{Cam}_{depth}$, resulting in an updated query feature $q_{i}^{updated}$. The overall sampling process can be formulated as follows.
$q_{i}^{updated}=\sum_{i=1}^{13}3DFA\left(q_{i},\bar{m}^{samp}_{i},Vol^{Cam}_{depth}\right)$ (17)
After the 3D cross-attention process, $q_{i}^{updated}$ already aggregates sufficient depth-aware visual features and is passed to the GCWSFusion module for feature fusion. Meanwhile, the set of 3D sampling points $m^{samp}$ for each T-primitive is also passed on to the GCWSFusion module for semantic-aware LiDAR feature aggregation.
III-G2 Gate-Concatenation and Weight-Summation Fusion (GCWSFusion) Module
Figure 8: Gate-Concatenation and Weight-Summation Fusion Module. The left part exhibits the multi-modality weighted summation fusion part, and the right part exhibits the multi-modality gated concatenation fusion part. The outputs of the two major parts are further fused to produce the final fused feature for each T-Primitive.
Given a set of 3D sampling points $m^{samp}$ for each T-primitive, we apply bilinear interpolation on $Vol^{Lidar}_{sem}$ to aggregate a semantic-aware LiDAR feature for each T-primitive. Then, both the semantic-aware LiDAR feature and the depth-aware visual feature of each T-primitive are fed into the GCWSFusion module for late-stage feature fusion. The detailed implementation of GCWSFusion module is demonstrated in Fig. 8. It consists of two central parts, the weighted summation fusion part, as highlighted in the green rectangle on the left of Fig. 8, and the gated concatenation fusion part, as highlighted in the pink rectangle on the right of Fig. 8. Both parts take the same inputs, but perform different feature fusion strategies.
For the weighted summation fusion part, the semantic-aware LiDAR feature and the depth-aware visual feature are summed in elements first, then passed through several linear layers followed by a softmax layer, resulting in two weight values, $W_{Lidar}$ and $W_{Vis}$. The $W_{Lidar}$ and $W_{Vis}$ are multiplied again with the corresponding semantic-aware LiDAR feature and depth-aware visual feature, respectively, resulting in the weighted LiDAR feature and the weighted visual feature. Lastly, the weighted LiDAR feature and the weighted visual feature are summed and passed through an extra set of linear layers, yielding the summation fusion feature. This design adaptively weights the dominance information across two sensors.
For the gated concatenation fusion part, the semantic-aware LiDAR feat and the depth-aware visual feat are passed through several linear layers, followed by a sigmoid function, respectively, resulting in a visual gate vector $G_{Vis}$ and a LiDAR gate vector $G_{Lidar}$. $G_{Vis}$ is element-wise multiplied with the semantic-aware LiDAR feature, followed by a skip connection that results in the gated LiDAR feature. $G_{Lidar}$ is element-wise multiplied with the depth-aware visual feature, followed by a skip connection that results in the gated visual feature. The gated LiDAR and visual features are concatenated along the feature channel dimension, followed by another set of linear layers that reduces the fused feature channel dimension, yielding the concatenation fusion feature. This design can dynamically amplify or suppress feature information within the two different modalities, thereby achieving denoising and information enhancement during the feature fusion process.
Lastly, the summation and concatenation fusion features are combined via feature-channel concatenation, followed by extra linear layers, yielding the final fused feature.
III-G3 3D Sparse Conv-based Self-Attention Module
Each T-primitive center $m_{i}$ is treated as a 3D point, and we can acquire a set of 3D points based on a set of T-primitives. Then, this T-primitive-based sparse 3D point cloud undergoes a coordinate transformation from Cartesian to Cylindrical coordinates, followed by voxelization, yielding a sparse voxel volume in cylindrical coordinates. Lastly, the 3D sparse convolution applies to this sparse voxel volume to perform the self-attention operation. Since our T-primitives are initialized in cylindrical coordinates, we perform the self-attention operation in the same coordinates to better preserve the scene’s fine-grained 3D geometry.
III-G4 T-Primitive Property Refinement
As presented in (18), the physical attributes of the three proposed T-Primitives are refined by processing each associated query vector through an MLP to produce updated attributes, which are subsequently added to the original attributes for refinement.
$(\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{c}_{i})=MLP(Q_{P_{i}})$
$(\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{c}_{i})=MLP(Q_{SQ_{i}})$
$(\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\left\{\hat{\omega_{i,k}}\right\}_{k=1}^{24},\hat{c}_{i})=MLP(Q_{SQIW_{i}})$
$\hat{P^{T}_{i}}=(m_{i}+\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{c}_{i})$
$\hat{P^{T}_{SQ_{i}}}=(m_{i}+\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{c}_{i})$
$\hat{P^{T}_{SQIW_{i}}}=(m_{i}+\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\left\{\hat{\omega_{i,k}}\right\}_{k=1}^{24},\hat{c}_{i})$ (18)
IV Experimental Results
IV-A Implementation Details
The TFusionOcc model takes six surround-view images and 10 LiDAR sweeps per data sample and leverages ResNet50-DCN [5] as a 2D Backbone, initialized with weights from FCOS3D [29]. The transformer layer is stacked four times iteratively. AdamW optimizer is used, with an initial learning rate of 2e-4 and a weight decay of 0.01. The learning rate is decayed using a multistep scheduler. The predicted occupancy has a resolution of $200\times 200\times 16$ for full-scale evaluation on SurroundOcc-nuScenes, Occ3D-nuScenes and nuScenes-C datasets. For data augmentation, we employ photometric distortions and a grid mask to the input images during training. Model training is conducted on two H100 GPUs with 94GB of memory.
IV-B Loss Function
To train TFusionOcc, we take advantage of the Lovász-Softmax loss [1] $L_{lovasz}$ and the binary cross-entropy loss $L_{BCE}$ to optimize the 3D semantic occupancy task. For depth supervision, a binary cross-entropy loss $L^{depth}_{BCE}$ is used to refine the depth distribution feature. The final loss is formulated as:
$L_{Total}=L_{lovasz}+\lambda\cdot L_{BCE}+L_{BCE}^{depth}$ (19)
where $\lambda$ balance the loss weight, in our case, we set $\lambda=10$. The final loss is applied only to the output of the last iterated transformer.
IV-C Dataset
The public nuScenes dataset [2], specifically designed for autonomous driving purposes, serves as the primary data source for our experiments. Furthermore, nuScenes-C [32], which not only augments the existing nuScenes dataset with additional simulated challenging scenarios, such as snow, fog, lowlight and brightness, etc, but also simulates several engineering failure cases, such as frame lost and camera crash, etc, is employed as a supplementary dataset to further evaluate the robustness of the proposed algorithm. Each corruption scenario in nuScenes-C has three severities, ranked easy, mid, and hard, respectively.
To perform the 3D semantic occupancy prediction task, we use dense labels obtained from SurroundOcc [31] and Occ3D [28], respectively. Since the test set lacks semantic labels, we train our model on the training set and evaluate its performance using the validation set. For SurroundOcc annotations, we set the range of the X and Y axes to [-50, 50] meters and the Z axis to [-5, 3] meters under LiDAR coordinates. For annotations from Occ3D, we set the range of the X and Y axes to [-40, 40] meters and the Z axis to [-1, 5.4] meters under the ego coordinates. The input images have a resolution of $1600\times 900$ pixels, while the final output of the semantic occupancy prediction branch is represented with a resolution of $200\times 200\times 16$. The annotations from SurroundOcc contain a total of 17 semantic classes, while the annotations from Occ3D comprise 18 semantic classes. Label 0 refers to free voxels in both annotations.
IV-D Model Performance Analysis
To demonstrate the performance of our proposed TFusionOcc, we evaluated it on the SurroundOcc-nuScenes and Occ3D-nuScenes benchmarks. For each benchmark, we evaluate our model across three primitive settings, including T-P (plain T-primitive), T-SQ (T-superquadric), and T-SQ-IW (deformable T-superquadric or T-superquadric with inverse-warp), with 12800 and 25600 total primitives, respectively. The results of the experiment on the SurroundOcc-nuScenes and Occ3D-nuScenes benchmarks are shown in Table II and Table III, respectively.
Method 2D-Backbone Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [3] R101-DCN C 23.96 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas [20] - C 28.66 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer [13] R101-DCN C 30.50 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer [8] R101-DCN C 30.86 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 C-CONet [30] R101 C 26.10 18.40 18.60 10.00 26.40 27.40 8.60 15.70 13.30 9.70 10.90 20.20 33.00 20.70 21.40 21.80 14.70 21.30 InverseMatrixVT3D [17] R101-DCN C 30.03 18.88 18.39 12.46 26.30 29.11 11.00 15.74 14.78 11.38 13.31 21.61 36.30 19.97 21.26 20.43 11.49 18.47 OccFormer [38] R101-DCN C 31.39 19.03 18.65 10.41 23.92 30.29 10.31 14.19 13.59 10.13 12.49 20.77 38.78 19.79 24.19 22.21 13.48 21.35 FB-Occ [14] R101 C 31.50 19.60 20.60 11.30 26.90 29.80 10.40 13.60 13.70 11.40 11.50 20.60 38.20 21.50 24.60 22.70 14.80 21.60 RenderOcc [22] R101 C 29.20 19.00 19.70 11.20 28.10 28.20 9.80 14.70 11.80 11.90 13.10 20.10 33.20 21.30 22.60 22.30 15.30 20.90 GaussianFormer [9] R101-DCN C 29.83 19.10 19.52 11.26 26.11 29.78 10.47 13.83 12.58 8.67 12.74 21.57 39.63 23.28 24.46 22.99 9.59 19.12 Co-Occ [21] R101 C 30.00 20.30 22.50 11.20 28.60 29.50 9.90 15.80 13.50 8.70 13.60 22.20 34.90 23.10 24.20 24.10 18.00 24.80 GaussianFormer2-256 [7] R101-DCN C 31.74 20.82 21.39 13.44 28.49 30.82 10.92 15.84 13.55 10.53 14.04 22.92 40.61 24.36 26.08 24.27 13.83 21.98 QuadricFormer [42] R101-DCN C 31.22 20.12 19.58 13.11 27.27 29.64 11.25 16.26 12.65 9.15 12.51 21.24 40.20 24.34 25.69 24.24 12.95 21.86 SurroundOcc [31] R101-DCN C 31.49 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 Inverse++ [19] R101-DCN C 31.73 20.91 20.90 13.27 28.40 31.37 11.90 17.76 15.39 13.49 13.32 23.19 39.37 22.85 25.27 23.68 13.43 20.98 LMSCNet [25] - L 36.60 14.90 13.10 4.50 14.70 22.10 12.60 4.20 7.20 7.10 12.20 11.50 26.30 14.30 21.10 15.20 18.50 34.20 L-CONet [30] - L 39.40 17.70 19.20 4.00 15.10 26.90 6.20 3.80 6.80 6.00 14.10 13.10 39.70 19.10 24.00 23.90 25.10 35.70 Co-Occ [21] - L 42.20 22.90 22.00 6.90 25.70 32.40 14.50 13.50 21.00 10.50 18.00 22.50 36.60 21.80 24.60 25.70 31.20 39.90 Co-Occ [21] R101-DCN C+L 41.10 27.10 28.10 16.10 34.00 37.20 17.00 21.60 20.80 15.90 21.90 28.70 42.30 25.40 29.10 28.60 28.20 38.00 GaussianFormer3D [39] R101-DCN C+L 43.30 27.10 26.90 15.80 32.70 36.10 18.60 21.70 24.10 13.00 21.30 29.00 40.60 23.70 27.30 28.20 32.60 42.30 OccFusion [18] R101-DCN C+R 33.97 20.73 20.46 13.98 27.99 31.52 13.68 18.45 15.79 13.05 13.94 23.84 37.85 19.60 22.41 21.20 16.16 21.81 OccFusion [18] R101-DCN C+L 43.53 27.55 25.15 19.87 34.75 36.21 20.03 23.11 25.25 17.50 22.70 30.06 39.47 23.26 25.68 27.57 29.54 40.60 OccFusion [18] R101-DCN C+L+R 43.96 28.27 28.32 20.95 35.06 36.84 20.33 26.22 25.86 19.17 21.27 30.64 40.08 23.75 25.56 27.63 29.82 40.82 GaussianFusionOcc [23] R101-DCN C+L 45.16 30.21 30.22 18.70 35.91 39.57 22.67 27.36 30.10 18.59 24.45 31.25 43.06 25.76 29.12 29.33 34.65 42.70 GaussianFusionOcc [23] R101-DCN C+L+R 45.20 30.37 30.43 18.54 36.23 39.66 22.57 27.35 30.30 19.14 24.56 31.95 42.60 25.82 29.48 29.70 34.78 42.95 OccCylindrical [16] Res-50 C+L 44.94 28.67 26.17 22.12 31.47 36.84 17.95 27.77 29.85 23.90 20.64 28.27 43.00 23.14 27.99 27.81 30.81 40.95 DAOcc [35] Res-50 C+L 45.00 30.50 30.80 19.50 34.00 38.80 25.00 27.70 29.90 22.50 23.20 31.60 41.00 25.90 29.40 29.90 35.20 43.50 TFusionOcc (T-SQ-12800) Res-50 C+L 46.70 30.76 30.67 21.27 33.12 38.96 22.09 27.54 30.47 22.87 21.65 29.86 42.63 27.18 29.80 30.91 36.81 46.26 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 46.92 31.21 31.65 20.13 34.75 39.70 23.08 27.00 30.89 23.33 22.02 30.40 43.11 27.78 30.25 31.47 37.15 46.60 TFusionOcc (T-P-25600) Res-50 C+L 45.49 30.04 30.15 19.58 34.20 38.64 22.10 26.60 29.85 22.24 22.44 29.65 41.08 25.19 28.43 29.71 35.77 45.01 TFusionOcc (T-SQ-25600) Res-50 C+L 46.14 30.88 30.87 21.26 33.68 39.25 22.54 28.34 31.45 23.61 22.80 30.16 41.76 26.12 29.23 30.26 36.77 46.06 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 47.01 31.47 31.96 20.78 34.57 39.73 24.02 27.82 31.19 23.82 23.05 30.67 43.17 27.18 30.13 31.18 37.52 46.71
TABLE II: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation set. Notion of modality: Camera (C), LiDAR (L), Radar (R). The best and second-best performance is indicated in bold red and bold blue, respectively.
Method Mask Backbone Modality mIoU $\bullet$ others $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [3] Res-101 C 6.06 1.75 7.23 4.26 4.93 9.38 5.67 3.98 3.01 5.90 4.45 7.17 14.91 6.32 7.92 7.43 1.01 7.65 BEVDet [6] Res-101 C 11.73 2.09 15.29 0.0 4.18 12.97 1.35 0.0 0.43 0.13 6.59 6.66 52.72 19.04 26.45 21.78 14.51 15.26 BEVFormer [13] Res-101 C 23.67 5.03 38.79 9.98 34.41 41.09 13.24 16.50 18.15 17.83 18.66 27.70 48.95 27.73 29.08 25.38 15.41 14.46 BEVStereo [12] Res-101 C 24.51 5.73 38.41 7.88 38.70 41.20 17.56 17.33 14.69 10.31 16.84 29.62 54.08 28.92 32.68 26.54 18.74 17.49 TPVFormer [8] Res-101 C 28.34 6.67 39.20 14.24 41.54 46.98 19.21 22.64 17.87 14.54 30.20 35.51 56.18 33.65 35.69 31.61 19.97 16.12 OccFormer [38] Res-101 C 21.93 5.94 30.29 12.32 34.40 39.17 14.44 16.45 17.22 9.27 13.90 26.36 50.99 30.96 34.66 22.73 6.76 6.97 RenderOcc [22] Swin-B C 26.11 4.84 31.72 10.72 27.67 26.45 13.87 18.20 17.67 17.84 21.19 23.25 63.20 36.42 46.21 44.26 19.58 20.72 CTF-Occ [28] Res-101 C 28.53 8.09 39.33 20.56 38.29 42.24 16.93 24.52 22.72 21.05 22.98 31.11 53.33 33.84 37.98 33.23 20.79 18.00 Inverse++ [19] Res-101 C 31.04 9.56 41.91 23.53 42.38 46.35 18.61 28.03 26.61 24.77 25.93 34.81 60.10 33.23 37.62 34.83 19.20 20.26 RadOcc [36] Swin-B C+L 49.38 10.90 58.20 25.00 57.90 62.90 34.00 33.50 50.10 32.10 48.90 52.10 82.90 42.70 55.30 58.30 68.60 66.00 OccFusion [18] R101-DCN C+L+R 46.67 12.40 50.30 31.50 57.60 58.80 34.00 41.00 47.20 29.70 42.00 48.00 78.40 35.70 47.30 52.70 63.50 63.30 OccFusion [18] R101-DCN C+L 48.74 12.40 51.80 33.00 54.60 57.70 34.00 43.00 48.40 35.50 41.20 48.60 83.00 44.70 57.10 60.00 62.50 61.30 GaussianFormer3D [39] R101-DCN C+L 46.40 9.80 50.00 31.30 54.00 59.40 28.10 36.20 46.20 26.70 40.20 49.70 79.10 37.30 49.00 55.00 69.10 67.60 RM2Occ [24] R101-DCN C+L 47.82 13.34 54.53 33.81 58.30 59.97 34.45 34.89 29.43 31.53 39.91 42.44 96.01 49.76 61.28 67.62 53.71 51.96 SDGOcc [4] Res-50 C+L 51.66 13.20 57.80 24.30 60.30 64.30 36.20 39.40 52.40 35.80 50.90 53.70 84.60 47.50 58.00 61.60 70.70 67.70 EFFOcc [26] Res-50 C+L 52.82 12.09 59.67 33.39 61.76 64.98 35.46 46.01 57.09 41.04 47.87 54.59 82.76 43.95 56.37 60.23 71.12 69.60 TFusionOcc (T-SQ-12800) Res-50 C+L 52.34 14.82 58.85 39.48 55.10 62.24 35.42 44.19 58.61 47.17 44.92 48.43 82.54 42.87 54.65 59.52 71.05 69.96 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 52.97 14.90 60.59 40.93 56.48 63.68 35.25 45.96 58.37 48.62 44.54 49.84 83.42 43.85 55.09 59.67 69.86 69.51 TFusionOcc (T-P-25600) Res-50 C+L 51.51 14.75 58.10 37.73 51.21 62.03 35.55 44.82 55.18 43.70 44.33 49.24 82.71 42.59 54.83 59.27 70.21 69.41 TFusionOcc (T-SQ-25600) Res-50 C+L 51.70 14.89 59.19 37.64 53.96 62.40 33.88 44.74 58.41 47.17 41.08 47.23 82.46 43.25 54.80 59.04 70.34 68.39 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 53.35 14.66 60.25 40.80 58.39 63.67 34.81 45.97 59.04 48.72 45.32 50.43 83.21 44.70 55.55 60.22 71.41 69.75
TABLE III: 3D semantic occupancy prediction results on Occ3D-nuScenes benchmark. Notion of modality: Camera (C), LiDAR (L), Radar (R). The best and second-best performance is indicated in bold red and bold blue, respectively.
In both benchmarks, TFusionOcc (T-SQ-IW) settings achieve the best and second-best performance, demonstrating the importance of the primitives’ shape variety and robustness. Specifically, our proposed method excels at detecting background objects, including drivable surfaces, other flat areas, sidewalks, terrain, man-made structures, and vegetation. In addition, our model performs significantly well on relatively small foreground objects, such as barriers, bicycles, cars, motorcycles, pedestrians, and traffic cones. Notably, our method, which only adopts a single 3D supervision signal, achieved par-performance in comparison with algorithms that leverage multi-task learning to receive dual supervision signals, such as DAOcc [35] in the SurroundOcc-nuScenes benchmark that incorporates an auxiliary 3D object detection head to leverage an extra 3D supervision signal during training, and SDGOcc [4] in the Occ3D-nuScenes benchmark, which incorporates an auxiliary 2D semantic segmentation head to utilize an extra 2D supervision signal during training. It is also worth noting that, as highlighted in the green and red rows in both benchmarks, for the same total number of primitive settings, the primitive with greater shape variety achieves better performance. In addition, increasing the number of primitives from 12,800 to 25,600 usually yields further gains, though the improvement is moderate compared with that from stronger primitive geometry. This suggests that primitive shape flexibility is a more decisive factor than simply increasing primitive count.
To further examine the performance and robustness of the proposed algorithm, we follow [18] to evaluate TFusionOcc in settings under challenging rainy and night-time scenarios, and the experiment results are shown in Table IV.
Method Backbone Modality Rainy IoU / mIoU Night IoU / mIoU Co-Occ [21] R101 C+L 40.30 / 26.60 35.60 / 14.60 OccFusion [18] R101-DCN C+L+R 42.67 / 27.39 41.01 / 16.61 GaussianFusionOcc [23] R101-DCN C+L 44.28 / 29.19 42.78 / 18.66 GaussianFusionOcc [23] R101-DCN C+L+R 44.36 / 29.86 42.51 / 18.45 OccCylindrical [16] Res-50 C+L 44.08 / 28.07 43.38 / 17.79 DAOcc [35] Res-50 C+L 44.51 / 29.65 42.85 / 18.53 TFusionOcc (T-SQ-12800) Res-50 C+L 45.79 / 30.28 44.74 / 18.81 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 45.81 / 30.74 44.82 / 19.07 TFusionOcc (T-P-25600) Res-50 C+L 44.77 / 29.22 43.25 / 17.83 TFusionOcc (T-SQ-25600) Res-50 C+L 45.40 / 30.38 43.86 / 18.95 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 46.25 / 30.82 44.97 / 19.67
TABLE IV: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation rainy and nighttime scenario subset. The best and second-best performance is indicated in bold red and bold blue, respectively.
In both challenging scenarios, our proposed method under the T-SQ-IW setting achieved the best and second-best performance, demonstrating its robustness and the importance of the geometric shape representation capability of primitives.
IV-E Performance Analysis on nuScenes-C Dataset
To thoroughly assess the robustness of the proposed algorithm under extreme conditions, such as snow, fog, and sensor malfunctions, we use the nuScenes-C dataset and evaluate TFusionOcc across varying levels of corruption and severity. The results of the experiment for several camera corruption settings are shown in Table V.
Method Clean IoU / mIoU Cam Crash IoU / mIoU Frame Lost IoU / mIoU Color Quant IoU / mIoU Motion Blur IoU / mIoU Bright IoU / mIoU Low Light IoU / mIoU Fog IoU / mIoU Snow IoU / mIoU OccCylindrical 44.94 / 28.67 41.76 / 23.38 40.84 / 21.87 43.64 / 25.38 41.45 / 22.57 44.78 / 27.88 43.48 / 22.17 44.59 / 27.38 38.92 / 18.29 DAOcc 45.00 / 30.50 42.75 / 22.35 42.16 / 20.78 43.88 / 27.46 42.75 / 26.25 43.72 / 26.78 41.25 / 21.15 43.92 / 26.87 40.72 / 19.42 TFusionOcc (T-P-25600) 45.49 / 30.04 42.59 / 22.34 41.97 / 20.78 43.14 / 24.42 40.45 / 20.04 44.81 / 29.03 41.24 / 18.62 44.14 / 27.43 40.58 / 15.62 TFusionOcc (T-SQ-25600) 46.14 / 30.88 42.59 / 23.50 41.93 / 22.04 43.03 / 25.65 38.36 / 20.25 45.41 / 29.92 41.99 / 20.98 44.90 / 28.85 40.09 / 16.33 TFusionOcc (T-SQ-IW-25600) 47.01 / 31.47 42.61 / 24.39 41.77 / 23.03 43.97 / 26.14 38.58 / 21.06 46.31 / 30.58 42.19 / 21.67 45.95 / 29.69 40.37 / 17.45
TABLE V: 3D semantic occupancy prediction results on nuScenes-C Camera Corruption benchmark. The best and second-best performance is indicated in bold red and bold blue, respectively.
In the camera corruption scenario, our model shows strong robustness against camera crashes, frame lost, brightness, and fog, as measured by the IoU and mIoU metric. Nevertheless, the motion blur and snow scenarios cause significant performance degradation in our model. With around 5.04% to 8.43% and 4.91% to 6.64% degradation under the IoU metric, 10.00% to 10.63% and 14.02% to 14.55% degradation under the mIoU metric for the motion blur and snow corruption case, respectively. This phenomenon could stem from the multi-stage feature fusion design strategy. Since frame lost and cam crash, both corruption results in the complete loss of information from the camera side. Our feature fusion module at each stage can easily filter out unnecessary visual features and rely solely on LiDAR features for occupancy prediction. As for brightness, fog, and low-light corruption, these conditions primarily degrade camera-side information quality and can be compensated for on the LiDAR side through our multi-stage feature fusion design. However, motion blur corruption causes structured information loss via convolution, bringing in inaccurate visual features into the feature fusion stage, resulting in a low-quality fused feature. Similarly, snow corruption introduces additional salt-and-pepper noise into the multi-stage feature fusion pipeline, resulting in a low-quality fused feature. Nevertheless, we consistently observe that more expressive primitives achieve stronger robustness under most corruption settings, following the trend T-SQ-IW >> T-SQ >> T-P, which further highlights the importance of geometric shape flexibility.
IV-F Performance Analysis on Varying Distance
[…]
Figure 9: Performance at different sector ranges for the 3D semantic occupancy prediction task. (a) mIoU performance at different sector ranges on the whole SurroundOcc-nuScenes validation set, (b) IoU performance at different sector ranges on the whole SurroundOcc-nuScenes validation set, (c) mIoU performance at different sector ranges on the SurroundOcc-nuScenes validation rainy scenario subset, (d) IoU performance at different sector ranges on the SurroundOcc-nuScenes validation rainy scenario subset, (e) mIoU performance at different sector ranges on the SurroundOcc-nuScenes validation night scenario subset, and (f) IoU performance at different sector ranges on the SurroundOcc-nuScenes validation night scenario subset. Better viewed when zoomed in.
[…]
IV-G Qualitative Study
To demonstrate the effectiveness and robustness of our method, we compare TFusionOcc in three different T-Primitive settings, such as T-P-25600, T-SQ-25600 and T-SQ-IW-25600, against other SOTA algorithms in challenging daytime, rainy and nighttime scenarios, as shown in Fig. 10, with the primary difference area highlighted in each scenario.
Figure 10: Qualitative results compare against other SOTA algorithms for daytime, rainy, and nighttime scenarios displayed in the upper, middle, and bottom sections, respectively. Better viewed when zoomed in.
In the daytime scenario, as shown at the top of Fig. 10, in the orange square highlighted area, benefit from multi-task learning, which provides an additional supervision signal, DAOcc successfully detects distant pedestrians and achieves similar performance as TFusionOcc. However, in the dark-red-square-highlighted area, DAOcc fails to detect the distant vehicle, demonstrating the superior performance of our proposed method. In the rainy scenario, as shown in the middle of Fig. 10, in the purple square highlighted area, our proposed method not only successfully detects all dynamic objects, but also preserves their general contours well. In the highlighted dark-blue-square area, our approach achieves par performance with other SOTA algorithms for detecting challenging pedestrians. In the nighttime scenario, as shown at the bottom of Fig. 10, our proposed method successfully detects very distant dynamic objects in the highlighted dark-red-square area, demonstrating its superior long-range dynamic object perception capability.
IV-H Model Efficiency
Method Modality Latency (ms) ($\downarrow$) Memory (GB) ($\downarrow$) Params M-CONet [30] C+L 416.00 12.10 125.43M OccCylindrical [16] C+L 361.39 5.97 111.62M OccMamba-384 [10] C+L 448.30 12.1 111.08M OccMamba-128 [10] C+L 269.30 6.8 64.58M TFusionOcc (T-SQ-12800) C+L 278.46 3.68 95.62M TFusionOcc (T-SQ-IW-12800) C+L 281.47 3.88 95.96M TFusionOcc (T-P-25600) C+L 385.49 4.73 97.51M TFusionOcc (T-SQ-25600) C+L 388.05 4.13 97.59M TFusionOcc (T-SQ-IW-25600) C+L 391.06 4.04 98.24M
TABLE VI: Model efficiency comparison of different methods. The experiments are performed on a single RTX4090 using six multi-camera images and 10 sweeps of LiDAR. For input image resolution, all methods adopt $1600\times 900$.
We evaluated the real-time performance of our model across various settings, comparing it with several other SOTA approaches on an RTX 4090 GPU, and presented the results in Table VI. The experiment results show the trade-off among the total number of primitives, memory consumption, and model inference speed. More primitives require more rendering operations and memory, leading to higher latency but better performance. Based on the main performance benchmark shown in Table II, the T-SQ-IW-12800 setting delivers the best balance between prediction and real-time performance.
IV-I Ablation Study
IV-I1 Ablation Study on Different Distribution and Primitives
Prob Distribution Query Number IoU ($\uparrow$) Gain ($\uparrow$) mIoU ($\uparrow$) Gain ($\uparrow$) T-P / G-P 25600 44.77 / 44.71 – / – 29.22 / 29.20 – / – T-SQ / G-SQ 25600 45.40 / 44.51 +0.63 / -0.20 30.38 / 30.25 +1.16 / +1.05 T-SQ-IW / G-SQ-IW 25600 46.23 / 45.58 +0.83 / +1.07 30.77 / 30.66 +0.39 / +0.41
TABLE VII: Ablation study on different probability distribution and primitives. Gain denotes the vertical performance change within each series (T-series / G-series). T refers to the Student’s t-distribution kernel, and G refers to the Gaussian kernel.
Table VII presents the ablation study results for different probability distributions and primitive settings. For each row in the table, for the same type of primitive and the same total number of primitives, with different distribution settings, the model obtained better performance, benefiting from the robustness property of the Student’s t-distribution. Notably, the geometric prediction capability benefits more than the semantic prediction capability. What’s more, by examining the table vertically, we can see that, regardless of the probability distribution kernel setting, the primitive with greater shape variety performs better, demonstrating the importance of the primitive’s geometric shape representation capability.
IV-I2 Ablation Study on Basis Field of T-SQ-IW Primitive
Basis field IoU ($\uparrow$) mIoU ($\uparrow$) $B_{1}\sim B_{24}$ 46.23 (Baseline) 30.77 (Baseline) $B_{10}\sim B_{24}$ 32.98 (-13.25%) 23.17 (-7.60%) $B_{1}\sim B_{18}$ 36.20 (-10.03%) 24.11 (-6.66%) $B_{1}\sim B_{9}$ 21.83 (-24.40%) 11.60 (-19.17%) No Inverse-Warp 20.24 (-25.99%) 13.87 (-16.90%)
TABLE VIII: Ablation study on basis field in T-SQ-IW primitive. $B_{10}\sim B_{24}$: primitive loses basic constant, linear, shear and taper deformation capability. $B_{1}\sim B_{18}$: primitive loses quadratic, radial bulge and smooth corner deformation capability. $B_{1}\sim B_{9}$: primitive further loses shear, twist and bending deformation capability.
Table VIII presents the results of the ablation study for each set of basis fields in the primitive T-SQ-IW. Since $B_{19}\sim B_{24}$ provides the highest level deformation, including quadratics, radial bulge, and smooth corner deformation, losing those basis fields causes 10.04% IoU and 6.66% mIoU degradation. In addition, because $B_{10}\sim B_{18}$ provides the second-highest level of deformation, including full shear, twist, and bending, losing $B_{10}\sim B_{24}$ results in performance drop. Surprisingly, the performance drop after losing all basis fields is similar to the loss $B_{10}\sim B_{24}$. This result is very likely because the remaining $B_{1}\sim B_{9}$ only provide basic constant, linear, shears, and taper deformation, which leads to limited improvement in the shape variety of the primitive. This result also demonstrates the importance of the high-level deformation basis field, which significantly improved the shape variety of the primitive, resulting in better overall 3D occupancy prediction performance.
IV-I3 Ablation Study on Skeleton Merge Module
Main-Skele Augment Overlap Remove Range Filter IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 46.20 (-0.04%) 30.70 (-0.09%) 46.07 (-0.17%) 30.76 (-0.03%) 46.08 (-0.16%) 30.69 (-0.10%) 41.96 (-4.28%) 29.92 (-0.87%) 29.35 (-16.89%) 22.31 (-8.48%)
TABLE IX: Ablation study on the Skeleton Merge Module. Main-Skele: LiDAR-side main skeleton branch. Augment: Camera-side augment skeleton branch.
Range Filter $r$ IoU ($\uparrow$) mIoU ($\uparrow$) 5 m 46.24 (Baseline) 30.79 (Baseline) 4 m 46.21 (-0.03%) 30.75 (-0.04%) 3 m 46.09 (-0.15%) 30.76 (-0.03%) 2 m 45.72 (-0.52%) 30.69 (-0.10%) 1 m 44.64 (-1.60%) 30.32 (-0.47%) 0.5 m 43.60 (-2.64%) 30.04 (-0.75%)
TABLE X: Ablation study on range filter hyperparameter $r$.
We conducted an ablation study for each component of the skeleton merge module and presented the results in Table IX. The experiment results prove the significant impact of the main skeleton structure, which supports the fundamental 29.92% mIoU and 41.96% IoU performance, and the auxiliary impact of the augmented skeleton, which further boosts 0.87% mIoU and 4.28% IoU performance. The main skeleton, based on LiDAR, provides precise 3D geometric information about the surrounding scene, especially the front faces of objects, and the augmented skeleton, based on the camera, provides additional information to mitigate the partial occlusion. Furthermore, we studied the impact of the range filter hyperparameter $r$ on the model’s overall performance and present the results in Table X. The results show that a smaller filter range forces the augmented skeleton to get too close to the main skeleton, preventing it from overcoming the LiDAR sensor’s occlusion limitation, resulting in performance degradation.
IV-I4 Ablation Study on Fused Depth Map Guided 3D Deformable Attention Module
Vis-Depth Fusion Depth-Fusion Vis-Depth LiDAR-Depth IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 46.06 (-0.18%) 30.57 (-0.22%) 39.80 (-6.44%) 25.28 (-5.51%) 44.01 (-2.23%) 28.98 (-1.81%) 43.99 (-2.25%) 28.93 (-1.86%) 40.33 (-5.91%) 24.71 (-6.08%)
TABLE XI: Ablation study on Depth-Fusion-Based 3D Deformable Attention Module. Vis-Depth Fusion: add the fused depth maps back to the visual features, resulting in depth-aware visual features. Depth-Fusion: fuse the visual depth maps with the LiDAR depth maps. Vis-Depth: visual depth maps. LiDAR-Depth: LiDAR depth maps.
We conducted an ablation study for each component of the fused-depth-guided 3D deformable attention module and presented the results in Table XI. In practice, Vis-Depth provides dense but less accurate depth maps that are good at capturing the partial occlusion of object structure. In contrast, LiDAR-Depth provides sparse but more accurate depth maps, especially good at capturing the front faces of objects in the scene. The experiment shows that Vis-Depth maps contribute significantly to the performance, even though they are less accurate. Relying solely on sparse but accurate LiDAR-Depth maps results in a performance degradation of 6.44% IoU and 5.51% mIoU, respectively. Furthermore, adding fused depth maps back to the original visual features results in depth-aware visual features contributing the second-largest performance improvement. The experiment results demonstrate the importance of depth map density and the necessity of multi-source depth map fusion and vis-depth feature fusion to boost model performance.
IV-I5 Ablation Study on GCWSFusion Module
We conducted an ablation study for each major component of the GCWSFusion module and present the experiment results in Table XII.
Gate-atten Weight-sum Skip-Connect WSFusion GCFusion IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 45.52 (-0.72%) 30.68 (-0.11%) 41.34 (-4.90%) 27.12 (-3.67%) 38.60 (-7.64%) 18.18 (-12.61%) 39.39 (-6.85%) 18.57 (-12.22%) 24.20 (-22.04%) 19.91 (-10.88%)
TABLE XII: Ablation study on the GCWSFusion Module. Skip-Connet: skip-connection structure at the last feature fusion stage. WSFusion: weighted summation fusion part. GCFusion: gated concatenation fusion part. Gate-atten: Gate attention mechanism in GCFusion module. Weight-sum: Weighted summation mechanism in the WSFusion module.
The results of the experiment show that the weight-summation mechanism plays a relatively more important role than the gate-attention mechanism in the overall GCWSFusion module. However, the WSFusion and GCFusion components contribute almost equally to the model’s final performance, with WSFusion slightly more important. Missing each part will cause 12.22% to 12.61% mIoU and 6.85% to 7.64% IoU performance degradation. The skip-connection structure also plays an important role during feature fusion; without it, the mIoU and IoU degrade by 22.04% and 10.88%, respectively.
V Conclusion
This paper presents TFusionOcc, a T-primitive-based object-centric multi-sensor fusion framework for 3D semantic occupancy prediction. The proposed method is built upon a family of Student’s t-distribution-based T-primitives, among which the deformable T-Superquadric with inverse warping serves as the key geometry-enhancing primitive. To further improve robustness, we introduce a unified probabilistic formulation based on Student’s t-distribution and the T-mixture model for joint occupancy and semantic modeling, together with a tightly coupled multi-stage fusion architecture for effective camera–LiDAR integration. Extensive experiments on SurrounOcc-nuScenes, Occ3D-nuScenes, and nuScenes-C demonstrate that TFusionOcc achieves strong performance and robustness in the most challenging conditions. In future work, we plan to further improve long-range perception and robustness under structured visual degradations, such as motion blur and snow.